# An Improved Two-Step Attack on CRYSTALS-Kyber

### v1

@@ 5. Conclusion
In this paper, we propose an efficient two-step attack including an enhanced CPA attack and a trail-and-error lattice attack for Kyber. In the CPA step, we modify the power leakage model to be more suitable for ARM Cortex-M4 architecture and further filter the candidate keys from the PCC results by using the Kendall’s rank correlation coefficient. The accuracy of finding the correct key is significantly improved. In the lattice step, we construct a trail-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time. Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads. Because the $basemul$ function called by different security levels of Kyber is the same, our work can be directly applied to the other parameters. Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grant 62104097, in part by the Key Research Plan of Jiangsu Province of China under Grant BE2022098, and in part by the Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001.

### v2

@@ 5. Conclusion
In this brief, we propose an efficient two-step attack including an enhanced CPA and a trial-and-error lattice attack for Kyber. In traditional CPA, false positives are the main factor affecting coefficients of secret key extraction in experiments (yang2023chosen, )and (kuo2023lattice, ). However, by employing the comprehensive CPA step we proposed, we can cleverly avoid false positives while reducing the number of required power traces in the practical attack. In the lattice step, we construct a trial-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time. Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads. Because the $basemul$ function called by different security levels of Kyber is the same, our work can be directly applied to the other parameters. Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.